OpenStudy (anonymous):

For what values of t, is tan t=cot t ? A. t= pi/4 + pi/2 x k where k is an integer B. no matter what value of t is, tan t is always different than cot t C. t=1 D. t= pi/4 + pi x k where k is an integer ***I don't really remember the rules.. could you please refresh me and explain? Thank youu :)

OpenStudy (mathstudent55):

\(\tan t = \dfrac{\sin t}{\cos t} \) \( \cot t = \dfrac{\cos t}{\sin t} \) Make these substitutions in your equation. Then cross multiply. What do you get?

OpenStudy (anonymous):

umm i'm not quite sure what you mean :(

OpenStudy (anonymous):

which equation am i substituting it in? :/

OpenStudy (anonymous):

sorry i'm a bit confused :(

OpenStudy (mathstudent55):

Your eq. has tan t and cot t. I wrote above two identities of what tan t is equal to and what cot t is equal to. Start with your equation, tan t = cot t. Now substitute tan t with what tan t is equal to. Substitute what cot t with what cot t is equal to. What do you get?

OpenStudy (anonymous):

ohhh so i cross multiply and get sin t^2 = cos t^2 ? :/

OpenStudy (mathstudent55):

\( \dfrac{\sin t}{\cos t} = \dfrac{\cos t}{\sin t} \) \(\sin^2 t = \cos^2 t\)

OpenStudy (anonymous):

ohhh okay what happens from here?

OpenStudy (mathstudent55):

\(\sin^2 t - \cos^2 t = 0\) \((\sin t + \cos t)(\sin t - \cos t) = 0\) \(\sin t = - \cos t\) or \( \sin t = \cos t\) Where do the sine and cosine functions have the same value?

OpenStudy (mathstudent55):